A 4.4 kg block 2.7 m off the ground sits at the top of a ramp. It is held and let go at the top of the ramp until it reaches the bottom where it is traveling 5.0 m/s. What is the power of friction on the block if it took the block 2.8 s to make it to the bottom of the ramp


1
Expert's answer
2021-06-29T08:17:49-0400
"P_f=m\frac{gh-0.5v^2}{t}\\
P_f=4.4\frac{(2.7)(9.8)-0.5(5)^2}{2.8}=22\ W"